OpenStudy (anonymous):

complete the square to find the center and radius of the circle whose equation is x^2 + y^2 -4y = 36

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

help meee

OpenStudy (dan815):

okay!! complete the square!

OpenStudy (dan815):

complete the square for the y part x^2 + y^2 -4y = 36 x^2+(y+a)^2 -a^2 = 36 what does a have to be?

OpenStudy (dan815):

x^2 + y^2 -4y = 36 x^2+(y+a)^2 -a^2 = 36; if i expand this out i get... x^2+ y^2+2ay +a^2-a^2=36 http://prntscr.com/700a09 , as the rest is all equal y^2+2ay +a^2-a^2 = y^2 -4y a=-2

OpenStudy (dan815):

questions concerns??!

OpenStudy (anonymous):

I think I got it. Thanks!